Question 4 (1 point) When rolling a die 114 times, what is the probability of rolling a 6 greater than 20 times? 1) 0.6553 2) 0.0948 3) 0.4395 4) 0.0474 5) 0.3447
Answer #1

Answer

We know that the probability of getting a 6 on a die roll = (favorable)/(total outcome) = 1/6 = 0.1667

sample size given in the question is n = 114

So, first we will find the mean and standard deviation value

Mean = n*p

where n = 114 and p = 0.1667

setting the given values, we get

Mean = mu = 114*0.1667 = 19

And standard deviation = k T Vp*(1 -p)*n

setting the values, we get

standard deviation = VO. 1 667 * (1-0.1667) * 114 1 5.8359 3.98

Now, using the formula P(x>X)= P(z>(ar{x}-mu)/sigma)

we have ar{x}=20, mu=19, sigma=3.98

setting the given values, we get

P(x>20)= P(z>(20-19)/3.98) = P(z>0.2513)

using the identity P(z>a) = 1-P(z<a)

we can write it as

P(z>0.2513) = 1-P(z<0.2513)= 1-0.6553 = 0.3447

We have used z distribution table to get the probability value.

So, required probability is 0.3447

option (5) is correct answer.